OpenStudy (anonymous):

evaluate without using a calculator cot^-1(1)

OpenStudy (abb0t):

remember, cot(x) = \(\frac{cos}{sin}\)

OpenStudy (anonymous):

does it matter which part i put for the inverse? Like can I put inverse cos/sin or cos/inverse sin?

OpenStudy (anonymous):

inverse of cot is just 1/cot

OpenStudy (anonymous):

or sin/ cos

OpenStudy (anonymous):

but 1/cot is easier for you here

OpenStudy (anonymous):

is this in radians?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

How would I find the answer without using a calculator? Inverse of cot(1)

OpenStudy (anonymous):

\[ \cot^{-1}(1) =x \]We can takes the \(\cot\) of both sides:\[ \cot(x) = 1 \]Then we can remember \( \cot(x) = \cos(x)/\sin(x) \)\[ \frac{\cos(x)}{\sin(x)} = 1 \]Finally we get\[ \cos(x) = \sin(x) \]

OpenStudy (anonymous):

i think inverse of cotangent(1) is also equal to inverse of cotangent (adjacent side/opposite side) of a right triangle which means the adjacent side and opposite side have the same lengths. In order to have the same lengths for a right triangle the angle should be 45 degrees since it is a right triangle

OpenStudy (anonymous):

Looking on the unit circle helps up imagine what this equation means: